what is scalar product...............................................?

Grade:12

3 Answers

SAI SARDAR
7 years ago
dear tushan,
                  Scalar product is defined as the product of the magnitudes of two vectors and the cosine of the angle between them.
Shreyas Patil
7 years ago
a scalar function of two vectors, equal to the product of their magnitudes and the cosine of the angle between them.
 
Kuldeep Pal
7 years ago
Scalar product is the Product of any Two vectors having angle between them (say ‘\alpha‘) with muliplication of cos\alpha.
Hence,
    \bar{a}.\bar{b} = abcos\alpha
   If,  \bar{a} = a1\hat{i} + a2\hat{j} and \tilde{b} = b1\hat{i} + b2\hat{j}
 Then ,   
             \bar{a}.\bar{b} = {a1b1 + a2b2}/{(a1)+ (a2)2)1/2.((b1)2 + (b2)2}